### v1

# Learning to Detect and Segment Mobile Objects from Unlabeled Videos

Abstract
Embodied agents must detect and localize objects of interest, e.g. traffic participants for self-driving cars. Supervision in the form of bounding boxes for this task is extremely expensive. As such, prior work has looked at unsupervised object segmentation, but in the absence of annotated boxes, it is unclear how pixels must be grouped into objects and which objects are of interest. This results in over-/under-segmentation and irrelevant objects. Inspired both by the human visual system and by practical applications, we posit that the key missing cue is motion: objects of interest are typically mobile objects. We propose MOD -UV, a Mobile Object Detector learned from Unlabeled Videos only. We begin with pseudo-labels derived from motion segmentation, but introduce a novel training paradigm to progressively discover small objects and static-but-mobile objects that are missed by motion segmentation. As a result, though only learned from unlabeled videos, MOD -UV can detect and segment mobile objects from a single static image. Empirically, we achieve state-of-the-art performance in unsupervised mobile object detection on Waymo Open [44], nuScenes [4], and KITTI [17] Dataset without using any external data or supervised models. Code is publicly available at %definitionshttps://github.com/YihongSun/MOD-UV.
Keywords:
Unsupervised Mobile Object Object Detection Videos
1 Introduction
Embodied agents such as self-driving cars must detect and localize objects of interest such as traffic participants to operate safely and effectively. Today, building such a detector requires the expensive and laborious annotation of millions of boxes over thousands of images. This process is so expensive that the largest detection dataset is orders of magnitude smaller than classification datasets and has much fewer classes. The limited set of classes further runs the risk of missing important object categories (e.g. snowplows for self-driving cars).
These concerns have motivated research into unsupervised object detection techniques that automatically discover objects from unlabeled data [51, 52, 6]. Under the hood, these techniques use self-supervised features to segment unlabeled images and produce candidate object annotations which are then used to train a detector. However, while promising, these approaches often produce many uninteresting and irrelevant “objects” in cluttered scenes (e.g. roads) and over- or under-segment objects of interest (e.g. multiple detections partitioning a large bus and a single detection grouping a row of parked cars). These failures shouldn’t be surprising: after all, how can a completely unsupervised feature representation encode that one assortment of windows, doors and wheels belongs together as an object of interest, but is a separate object from other nearby windows, doors and wheels?
In this paper, we argue that a key missing cue is motion. As humans, we know to group the windows, doors and wheels together because we have seen this assortment move together, also known as Gestalt principle of common fate [34]. Having seen this assortment move together multiple times in the past, we can now group these correctly into an object even when the object (a car) is static. Indeed, motion-based grouping is one of the first forms of grouping to appear developmentally in human infants [43] and can potentially bootstrap other grouping cues [33]. Motion is important not just perceptually, but also practically: Robots performing navigation tasks must plan their trajectories carefully around objects that can move. Identifying mobile objects is also important in manipulation tasks because it provides a view of the available affordances.
With these motivations in mind, we propose an unsupervised approach to build a detector for mobile objects as shown in Figure 1. Our approach trains on unlabeled videos alone and produces a mobile object detector that can run on static frames. We build upon recent advances in unsupervised learning of depth and motion segmentation from unlabeled video to automatically produce pseudo training labels without supervision. However, such unsupervised motion segmentation poses two key challenges. One, it only identifies moving objects and does not include objects (e.g. parked cars) that are static but mobile. Two, it only identifies nearby objects, since the pixel motion of faraway objects is too subtle to discern. We propose a new training scheme to address these challenges, resulting in a final mobile object detector that detects more mobile objects than the initial motion segmentation.
We test MOD -UV on self-driving scenes but evaluate on a variety of datasets. Specifically, we compare to recent state-of-the-art unsupervised object detectors and demonstrate improvements across the board, with notable improvements in Box AR by 6.6 on Waymo Open [44], 4.9 on nuScenes [4] and 6.2 on KITTI[17].
In sum, our contributions are:
1.
We argue that motion is an important cue for learning what makes an object.
2.
We propose a new training scheme that trains on unlabeled video to produce a mobile object detector that can run on static images.
3.
We demonstrate marked improvements over unsupervised object detection baselines across a range of datasets and metrics.
2 Related Work
2.0.1 Unsupervised Object Detection/Discovery from Images.
Learning to identify and localize objects from unlabeled images is a challenging task, as object information must be obtained without any explicit human annotations. A long line of work seeks to discover prominent objects in large image collections. Early work used probabilistic Hough transforms to match proposals across images[9]. This approach was improved upon by framing the problem as an optimization problem over a graph where nodes are object proposals and edges are visual similarities[48] and scaled up to larger collections [49, 50]. However, this approach is fundamentally limited by the quality of the object proposals. More recently, Locatello et al. [30] and DINOSAUR [36] consider object discovery as object-centric learning [18] and leverage Slot Attention to decompose a complex scene into independent objects. However the reconstruction objective is difficult to scale and is likely to discover irrelevant objects as well.
More recent work has relied on the fact that bottom-up segmentation algorithms when applied to self-supervised pretrained representations yield good object proposals. LOST [39], FOUND [40] and FreeSOLO [51] generate pseudo mask labels from DINO [7] features to train their downstream detectors. Similarly, MaskDistill [47] distills a single class-agnostic initial mask from the affinity graph produced by DINO [7]. TokenCut [53] and CutLER [52] use Normalized Cuts [37] instead to generate pseudo-labels. HASSOD [6] leverages hierarchical adaptive clustering to generate pseudo-labels, which improves the detection of small objects and object parts. This line of work now produces detectors that can run on static images, similar to our work. However, the detected objects can often be irrelevant (e.g. buildings and road) or over- or under-segment objects of interest. In contrast, our proposed MOD -UV discovers and detects a more meaningful and practical set of mobile objects instead, and can learn from their apparent motion in unlabeled videos only without relying on any additional datasets.
2.0.2 Unsupervised Object Detection/Discovery from 3D.
In addition to unlabeled images, 3D information can also be leveraged for discovering objects. Herbst et al. [21] reconstructs the same scene from multiple traversals via RGB-D inputs and identify discrepancies as discovered mobile objects. MODEST [57] extends upon this and builds an unsupervised 3D detector from traversals of LiDAR scans. Garcia et al. [16] discovers salient objects by late-fusing color and depth segmentation from RGB-D inputs. Tian et al. [46] generates candidate segments via 3D point cloud segmentation from LiDAR recordings. In comparison, MOD -UV does not require any additional sensors or modalities beyond unlabeled videos, which allows our method to work in more general settings.
2.0.3 Unsupervised Video Segmentation.
Another class of related work segments out the moving objects from a video without annotations. Here the input is a video and the output is a binary segmentation of the moving objects. Kwak et al. [25], Xie et al. [54], Lamdouar et al. [26], Yang et al. [56], Zhang et al. [58], Zhou et al. [59], and FODVid [41] perform this task by leveraging optical flow information from an input video. Additionally, Lian et al. [28] improves upon cases of articulated/deformable objects and shadow/reflections by relaxing the common fate assumption. Another line of work uses a reconstruction objective to identify the moving object. Du et al. [13] models explicit physical dynamics and reconstructs 3D object geometry by exploiting motion cues in videos. Bao et al. [2] improves training for object-centric representation via an additional motion segmentation regularization, while SAVi++ [14] incorporates depth values from LiDAR sensor when training an object-centric video model. Aydemir et al. [1] extends DINOSAUR [36] to video learning while proposing a masking strategy for tokens in the feature space for efficiency and regularization. Unlike MOD -UV, these approaches do not build a static image detector. However, the output segmentation can be used as an initialization for our approach.
Closer to our work, Croitoru et al. [11] and Choudhury et al. [10] train a single-frame binary segmentation network on frames of a video as input and leverage object motion as supervision. Furthermore, LOCATE [42] applies graph-cut to obtain binary motion mask from DINO [7] and optical flow feature similarities, which in turn is treated as pseudo-labels for bootstrapped self-training of a downstream segmentation network. However, these techniques can only detect a single salient object per frame. In contrast, MOD -UV generalizes to multi-object detection during inference beyond single-object saliency detection.
3 Method
Problem setup: We assume an uncurated collection of unlabeled videos as input. In particular, we assume that these videos are obtained by an embodied agent observing, and optionally acting in the world. Solely from the unlabeled videos, the goal is to learn a detector that operates from a single frame and can detect and segment all mobile objects that can move of their own volition.
3.1 Initialization with Unsupervised Motion Segmentation
A key insights in MOD -UV is that if an object can move, it is likely that it does move many times in the collected data. Thus, we start by identifying moving objects in the videos; they can be initial seeds for learning about mobile objects.
Fortunately, the task of identifying independently moving pixels from unlabeled videos is a well-studied one[3, 38, 55, 35]. In particular, many recent techniques have been proposed that learn motion segmentation without supervision from unlabeled videos. Many of these techniques also produce depth and camera motion [35, 31, 27, 23]. Here, we use our prior work, Dynamo-Depth [45]. Dynamo-Depth trains on unlabeled videos and learns both a monocular depth estimator as well as a motion segmentation network. We use the outputs of these trained networks on our unlabeled videos as a starting point. Concretely, we denote the input set of unlabeled videos as , with each video containing consecutive frames and known camera intrinsics. For each frame , we obtain its estimated motion mask and estimated monocular depth .
With the given binary motion mask , we first need to partition the moving pixels into instance-level labels. While disjoint moving regions can be easily separated, multiple moving objects in the same region would require additional information (e.g. 3D information) to separate. Therefore, for each image , we project the corresponding moving pixels in into pseudo 3D point clouds via the estimated monocular depth and inverse camera intrinsics .111 denotes the conversion to homogeneous coordinates
Then, we cluster via DBSCAN [15] to get a pseudo depth-aware partition of the motion mask , which we treat as the initial pseudo-labels, .
We evaluate the quality of these pseudo-labels qualitatively in Figure 2 and quantitatively in the top rows of Table 4. We find that these pseudo-labels have high precision, but have two severe limitations. First, they only identify moving objects, so they miss objects that are static but can move (e.g. parked cars) Second, they miss almost all faraway objects which tend to be small. This is because the apparent pixel motion of faraway objects is very hard to detect.
To tackle this issue of limited recall, we propose two self-training stages, Moving2Mobile and Large2Small, that progressively recover more mobile objects in the scene by aligning the training distribution of static and small objects with the available large moving objects in the initial pseudo-labels, respectively.
3.2 Self-Training for Unsupervised Mobile Object Detection.
Here, we describe each self-training stage of MOD -UV, as we progressively discover mobile objects to train the final mobile object detector.
3.2.1 Moving2Mobile: Learning to Detect Static Objects.
As shown in the left of Figure 2, the initial pseudo-labels , while having high precision, fail to capture the large static objects, e.g. the black sedan in the bottom. However, a parked black sedan looks the same as a moving black sedan if all one has is a single frame. In other words, moving and static objects are indistinguishable when observed from a single frame.
Thus, in Moving2Mobile, we simply train a detector to reproduce the pseudo-labeled instances in , but with only a single frame as input. Since object motion is not apparent in a single frame, this detector cannot distinguish moving objects from static ones and thus is forced to detect anything that share the appearance of moving objects, thus detecting static mobile objects as well.
However, there may exist domain-specific statistical regularities that give hints to object motion even in a single frame. For example, lit-up tail-lights might indicate that the car is stopped, while a highway background might suggest that the cars are moving. To prevent the detector from overfitting to these priors, we stop training early. Afterwards, we treat the high confidence predictions by the detector as the pseudo-labels for the next round .
3.2.2 Large2Small: Learning to Detect Small Objects.
As shown in Figure 2, , pseudo-labels after Moving2Mobile, appropriately recovers the large static objects, however, the smaller objects remain absent. Intuitively, faraway objects have much smaller apparent pixel motion (and thus are absent from ) and also look different from large moving objects (and thus are absent from ).
To learn to detect small objects, we create a new training dataset by scaling down both the image and the pseudo-labels (while also padding the image to maintain image size). We then train a separate “small object” detector by training on this new dataset. Intuitively, by training on the scaled down training pair, the output detector would need to detect the same object at a much lower scale, directly promoting the extension to small objects.
Also, since came from a heavily-regularized detector, we maintain and finetune the pseudo-labels for larger objects by training a second detector from scratch in parallel, on the training pair without down-scaling or padding. Notably, this is different from traditional scale jittering, since the singular detector would be discouraged from detecting small objects at larger scales due to the limitations in . Upon convergence, we have a Large-object detector trained at original scale and another Small-object detector trained at a reduced scale. After aggregating their predictions and resolving conflicting proposals222Details on the mask aggregation are found in Appendix 0.A.1., we obtain the final pseudo-labels, . We note that separating out large and small object detectors in this way has been explored in supervised face detection [22].
3.2.3 Final Round of Self-Training.
As shown in the right of Figure 2, , the final pseudo-labels after Large2Small, successfully recovers both static and small objects without introducing excessive false-positives. From here, we train the final detector from scratch, on the training pair to convergence.
3.3 Implementation details
We follow the official code release by Dynamo-Depth[45] and train the system on Waymo Open [44] only to obtain the motion segmentation network and monocular depth estimator. In the initial pseudo-label generation stage, we binarize the estimated motion mask via a threshold of and cluster the pseudo 3D points via DBSCAN [15] using a 10-by-10 local pixel neighborhood connectivity, Eps at 0.03, and required number of samples for core point at 100.
We adopt Mask R-CNN [19] with a ResNet-50[20] backbone as the detector architecture. We initialize the backbone via two strategies, namely MoCo v2[8] on randomly sampled Waymo [44] patches and MoCo v2 on ImageNet [12], denoted as MOD -UV‡ and MOD -UV, respectively. We use Adam optimizer [24] with initial learning rate of 1e-4 and decay by after 10 epochs.
During Moving2Mobile, we train a detector for 3 epochs, with scale jittering from to . Since early-stopping is applied, we adopt a lower confidence threshold of to compute the next round pseudo-labels . During Large2Small, we train both the large and small detectors for 20 epochs, with fixed scale jittering at and , respectively. As both are trained to convergence, we adopt a higher confidence threshold of and , respectively, to compute the next round pseudo-labels with aggregation. For Final round, we train the detector from scratch for 20 epochs, with scale jittering from to . The self-training in MOD -UV takes 27 hours on 1 NVIDIA A6000 GPU.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets.
For evaluation, we focus our attention to self-driving datasets since uncurated, unlabeled video data is available and detecting mobile objects is of interest for autonomous vehicles. We train both MOD -UV‡ and MOD -UV on Waymo [44] and compare our performance with baselines on Waymo. We then also evaluate generalization to nuScenes [4], KITTI [17], and COCO [29].
We split the 798 sequences from Waymo train set into 762 for training and 36 for validation. After method development concludes, we evaluate on the held-out 1,881 test images (averaging 28.4 mobile instances per image) from Waymo val set [32]. Additionally, nuScenes, KITTI, and COCO are only used for evaluating generalization. We test on 3,249 front-camera images (average of 8.2 mobile instances per image) from the nuImage validation set for nuScene and 7,481 images (average of 6.9 mobile instances per image) in the 2D Detection training set for KITTI. For COCO, we evaluate on 870 images (average of 3.8 mobile instances per image) in COCO val 2017 that contain ground vehicles.333Additional dataset information is found in Appendix 0.B.
4.1.2 Baselines.
For the task of unsupervised mobile object detection, there is no directly comparable baselines to the best of our knowledge. Therefore, we consider methods for unsupervised object detection, namely CutLER [52] and HASSOD [6], as the closest points of comparison.
CutLER [52] uses normalized cuts on DINO features (trained on ImageNet) to generate pseudo labels that are used to train a Cascade Mask R-CNN [5] with ResNet-50 backbone. We evaluate the official checkpoint.
HASSOD [6] is a follow-up to CutLER. It discovers objects on COCO [29] via a hierarchical adaptive clustering of DINO features (trained from ImageNet). The hierarchical clustering yields three “levels”: objects, object parts, and object sub-parts. For consistency, we consider all three hierarchical levels for evaluation. As before, we use its official released checkpoint. Since both MOD -UV‡ and MOD -UV are trained on Waymo [44], we also consider a version of HASSOD solely trained on Waymo Open [44], which we denote as HASSOD†.
CutLER[52] and HASSOD[6] are trained to detect all objects in an image regardless of their ability to move. However, our evaluation and approach is focused on mobile objects. We therefore also consider an “oracle” version of these baselines where we additionally remove any CutLER and HASSOD predictions that overlap by less than 0.1 in IoU with the ground truth instances. These oracles, namely CutLER∗, HASSOD∗, and HASSOD†∗, are grayed out in the tables to indicate additional ground-truth-based filtering.
We also consider a fully-supervised Mask R-CNN (Sup. Mask R-CNN) trained on COCO [29] for an oracle comparison, marked in gray.
4.1.3 Metrics.
We evaluate both Average Recall (AR and AR) and Average Precision (AP and AP). Since the task is unsupervised and no semantic information is given during training, we follow prior arts [52, 6] and evaluate class-agnostic AR and AP by treating all predicted and ground truth instances as a single class of “foreground” or “mobile” objects.444Complete benchmarks, repeated training performance, and additional ablation studies are found in Appendix 0.C.
Method | Train | Init. | Box | Mask | ||||||||
AR0.5 | AR | AR | AR | AR | AR0.5 | AR | AR | AR | AR | |||
Sup. Mask R-CNN [19] | 54.3 | 31.9 | 13.4 | 53.0 | 79.8 | 48.9 | 27.5 | 9.2 | 46.3 | 78.9 | ||
CutLER[52] | IN | IN | 20.9 | 11.7 | 1.3 | 17.3 | 54.1 | 20.0 | 10.7 | 0.9 | 14.9 | 52.7 |
HASSOD[6] | COCO | IN | 21.9 | 12.7 | 1.8 | 17.6 | 59.5 | 20.7 | 11.0 | 1.2 | 14.2 | 55.6 |
HASSOD†[6] | W | IN | 15.3 | 8.3 | 0.5 | 11.0 | 43.7 | 14.6 | 7.2 | 0.2 | 7.9 | 42.6 |
MOD -UV‡ (ours) | W | W | 40.0 | 17.5 | 8.4 | 25.7 | 46.4 | 35.4 | 14.6 | 6.8 | 21.2 | 40.4 |
MOD -UV (ours) | W | IN | 39.9 | 19.3 | 8.6 | 28.5 | 54.1 | 35.8 | 16.4 | 7.2 | 23.8 | 47.3 |
AP50 | AP | AP | AP | AP | AP50 | AP | AP | AP | AP | |||
Sup. Mask R-CNN [19] | 46.1 | 25.7 | 8.0 | 42.2 | 72.2 | 41.9 | 22.4 | 5.3 | 36.2 | 73.0 | ||
CutLER[52] | IN | IN | 8.8 | 5.0 | 0.5 | 3.9 | 32.0 | 9.1 | 5.2 | 0.0 | 3.4 | 34.6 |
HASSOD[6] | COCO | IN | 5.0 | 3.1 | 0.1 | 2.5 | 28.4 | 5.0 | 2.8 | 0.0 | 2.0 | 27.7 |
HASSOD†[6] | W | IN | 3.7 | 2.2 | 0.0 | 1.1 | 17.2 | 3.9 | 2.0 | 0.0 | 0.9 | 18.3 |
MOD -UV‡ (ours) | W | W | 26.1 | 10.9 | 4.2 | 16.2 | 32.0 | 25.0 | 9.5 | 3.7 | 14.1 | 28.7 |
MOD -UV (ours) | W | IN | 26.3 | 12.6 | 4.9 | 17.9 | 41.0 | 25.1 | 11.1 | 4.5 | 15.6 | 36.3 |
CutLER∗[52] | IN | IN | 14.3 | 7.4 | 1.3 | 10.7 | 38.4 | 14.3 | 7.4 | 0.9 | 9.6 | 41.0 |
HASSOD∗[6] | COCO | IN | 15.3 | 8.5 | 1.4 | 10.9 | 42.6 | 15.0 | 7.7 | 1.0 | 8.9 | 41.3 |
HASSOD†∗[6] | W | IN | 9.3 | 4.7 | 0.6 | 6.5 | 25.2 | 9.6 | 4.5 | 0.1 | 4.9 | 27.0 |
Method | Train | Init. | Box | Mask | ||||||||
AR0.5 | AR | AR | AR | AR | AR0.5 | AR | AR | AR | AR | |||
CutLER[52] | IN | IN | 28.9 | 16.0 | 3.8 | 23.4 | 56.1 | 27.8 | 14.3 | 3.0 | 20.4 | 53.0 |
HASSOD[6] | COCO | IN | 30.9 | 17.0 | 5.0 | 24.0 | 56.8 | 29.7 | 14.7 | 3.9 | 20.1 | 53.6 |
HASSOD†[6] | W | IN | 24.3 | 12.7 | 2.7 | 18.0 | 48.2 | 23.0 | 10.7 | 1.9 | 14.0 | 45.7 |
MOD -UV‡ (ours) | W | W | 42.1 | 17.3 | 8.7 | 24.2 | 39.8 | 36.4 | 13.9 | 8.2 | 18.1 | 29.7 |
MOD -UV (ours) | W | IN | 48.9 | 21.9 | 12.0 | 29.8 | 48.2 | 42.3 | 18.3 | 10.7 | 24.0 | 39.1 |
AP50 | AP | AP | AP | AP | AP50 | AP | AP | AP | AP | |||
CutLER[52] | IN | IN | 6.0 | 3.7 | 0.3 | 3.1 | 23.7 | 5.9 | 3.5 | 0.1 | 2.8 | 22.7 |
HASSOD[6] | COCO | IN | 3.9 | 2.2 | 0.1 | 2.1 | 20.5 | 3.8 | 2.0 | 0.1 | 1.8 | 19.5 |
HASSOD†[6] | W | IN | 3.6 | 2.2 | 0.0 | 1.7 | 15.2 | 3.6 | 1.8 | 0.0 | 0.9 | 14.6 |
MOD -UV‡ (ours) | W | W | 18.8 | 7.3 | 2.6 | 10.4 | 22.3 | 17.1 | 6.0 | 2.4 | 8.0 | 17.2 |
MOD -UV (ours) | W | IN | 23.6 | 10.7 | 4.3 | 14.9 | 31.5 | 21.8 | 9.0 | 3.8 | 12.2 | 25.6 |
CutLER∗[52] | IN | IN | 15.6 | 8.2 | 2.3 | 12.6 | 38.4 | 15.2 | 7.7 | 1.7 | 11.4 | 37.0 |
HASSOD∗[6] | COCO | IN | 18.6 | 9.5 | 2.3 | 13.3 | 38.2 | 17.9 | 8.6 | 1.8 | 11.5 | 36.2 |
HASSOD†∗[6] | W | IN | 13.0 | 6.3 | 1.2 | 9.1 | 27.9 | 12.8 | 5.6 | 1.1 | 7.5 | 27.2 |
Method | Train | Init. | KITTI [17] | COCO [29] | ||||||||
AR0.5 | AR | AR | AR | AR | AR0.5 | AR | AR | AR | AR | |||
CutLER[52] | IN | IN | 50.9 | 24.4 | 11.2 | 23.0 | 42.9 | 49.0 | 27.9 | 9.8 | 33.9 | 61.6 |
HASSOD[6] | COCO | IN | 52.4 | 26.4 | 13.4 | 23.9 | 47.1 | 51.2 | 27.9 | 12.5 | 36.2 | 51.7 |
HASSOD†[6] | W | IN | 49.9 | 23.7 | 15.0 | 22.1 | 37.4 | 48.8 | 26.1 | 12.5 | 31.2 | 50.7 |
MOD -UV‡ (ours) | W | W | 66.1 | 29.3 | 21.2 | 29.0 | 39.6 | 55.9 | 23.3 | 20.1 | 26.9 | 25.5 |
MOD -UV (ours) | W | IN | 68.1 | 32.6 | 23.7 | 32.0 | 44.3 | 58.3 | 26.6 | 22.3 | 29.8 | 32.0 |
AP50 | AP | AP | AP | AP | AP50 | AP | AP | AP | AP | |||
CutLER[52] | IN | IN | 18.6 | 8.6 | 0.4 | 5.6 | 24.0 | 9.8 | 5.6 | 0.4 | 2.4 | 21.6 |
HASSOD[6] | COCO | IN | 14.3 | 7.2 | 0.3 | 5.1 | 29.5 | 3.1 | 1.4 | 0.1 | 1.9 | 7.6 |
HASSOD†[6] | W | IN | 16.7 | 7.5 | 0.6 | 6.8 | 19.3 | 4.7 | 2.7 | 0.3 | 3.1 | 10.0 |
MOD -UV‡ (ours) | W | W | 38.5 | 15.8 | 9.8 | 15.3 | 26.2 | 14.1 | 5.8 | 5.1 | 7.6 | 5.9 |
MOD -UV (ours) | W | IN | 38.9 | 18.0 | 11.7 | 18.0 | 28.4 | 14.2 | 6.6 | 5.6 | 9.0 | 6.6 |
CutLER∗[52] | IN | IN | 28.4 | 12.3 | 3.5 | 9.9 | 28.8 | 24.3 | 12.8 | 4.3 | 12.1 | 40.1 |
HASSOD∗[6] | COCO | IN | 33.4 | 15.4 | 4.4 | 11.5 | 34.9 | 21.3 | 10.1 | 5.4 | 13.8 | 18.9 |
HASSOD†∗[6] | W | IN | 30.8 | 12.7 | 5.7 | 11.1 | 23.3 | 22.8 | 10.6 | 5.8 | 12.7 | 21.3 |
Pseudo | # Epochs | All | Static | Moving | |||||||||
Masks | in M2M | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR |
Contour | ✗ | 0.9 | 0.0 | 0.3 | 6.4 | 0.6 | 0.0 | 0.0 | 2.0 | 3.8 | 0.0 | 1.0 | 14.4 |
Depth | ✗ | 1.5 | 0.0 | 0.5 | 10.0 | 0.7 | 0.0 | 0.0 | 2.4 | 6.3 | 0.0 | 1.5 | 23.8 |
Depth | 1 | 4.4 | 0.0 | 3.2 | 25.9 | 7.1 | 0.0 | 2.2 | 21.1 | 11.4 | 0.0 | 6.7 | 34.7 |
Depth | 3 | 5.3 | 0.0 | 4.6 | 29.0 | 9.1 | 0.0 | 3.5 | 25.8 | 12.4 | 0.0 | 8.7 | 35.1 |
Depth | 5 | 4.3 | 0.0 | 3.2 | 24.7 | 6.7 | 0.0 | 1.8 | 20.2 | 11.1 | 0.0 | 7.1 | 32.9 |
Depth | 20 | 4.1 | 0.0 | 2.3 | 25.6 | 6.6 | 0.0 | 1.3 | 20.8 | 10.6 | 0.0 | 5.2 | 34.4 |
Stage 1 | Stage 2 | Stage 3 | Box AR | Box AP | ||||||||||
M2M | L2S | Final | AR0.5 | AR | AR | AR | AR | AP50 | AP | AP | AP | AP | ||
✗ | ✗ | ✓ | 18.0 | 7.8 | 0.3 | 9.7 | 43.3 | 10.5 | 4.2 | 0.4 | 3.5 | 25.3 | ||
✗ | only | ✗ | 12.9 | 5.6 | 0.0 | 4.5 | 38.3 | 7.6 | 2.9 | 0.4 | 1.1 | 22.3 | ||
✗ | only | ✗ | 23.0 | 10.2 | 3.0 | 18.3 | 28.3 | 11.8 | 5.1 | 1.6 | 8.9 | 15.0 | ||
✗ | ✓ | 28.4 | 12.6 | 2.3 | 20.8 | 47.7 | 16.4 | 7.0 | 1.6 | 10.6 | 32.8 | |||
✓ | ✗ | ✗ | 23.2 | 10.1 | 1.0 | 16.0 | 45.1 | 13.7 | 5.6 | 0.5 | 7.8 | 28.1 | ||
✓ | ✗ | ✓ | 28.1 | 12.5 | 2.2 | 20.3 | 48.4 | 16.7 | 7.1 | 1.4 | 10.8 | 33.5 | ||
✓ | only | ✗ | 21.3 | 9.4 | 0.2 | 14.9 | 45.2 | 12.8 | 5.7 | 0.5 | 7.2 | 31.2 | ||
✓ | only | ✗ | 31.0 | 13.4 | 5.8 | 22.4 | 32.5 | 17.1 | 7.2 | 2.9 | 13.6 | 15.0 | ||
✓ | ✓ | 40.0 | 17.5 | 8.4 | 25.7 | 46.4 | 26.1 | 10.9 | 4.2 | 16.2 | 32.0 |
4.2 Unsupervised Mobile Object Detection and Segmentation
4.2.1 In-Domain Performance on Waymo.
Table 1 compares MOD -UV‡ and MOD -UV against CutLER, HASSOD, and HASSOD† in unsupervised mobile object detection on Waymo. We also report performance for a MaskR-CNN trained on COCO as an oracle in the first row of Table 1.
Recall. We report AR0.5 (average recall with IoU), AR, AR, AR, and AR for both box and mask predictions. MOD -UV significantly outperforms prior arts across all recall metrics except the recall for large objects where it is comparable. The improvement is especially large for small objects (4.7 higher AR than the nearest competitor). Compared to a supervised Mask R-CNN trained on COCO [29], MOD -UV closes the gap in Box AR from 11.6 to 4.8. Our gains are also much larger on the AR0.5 metric (nearly 2 prior state-of-the-art). This suggests that we detect significantly more objects than prior work, but their localization can be improved. Even so, on overall AR we still show a -point improvement. Additionally, although HASSOD† is trained in-domain on Waymo, it underperforms HASSOD on Waymo, which we suspect is due to the uncurated nature of self-driving scenes.
Precision. We report AP at an overlap threshold of 0.5, as well as AP, AP, AP and AP. Since CutLER and HASSOD are trained to detect all objects in an image regardless of their ability to move, we also compare to oracle versions of these techniques with ground-truth-based filtering. On Waymo, MOD -UV significantly outperforms prior arts across all precision metrics. Even with ground-truth filtering (in gray), MOD -UV still consistently outperforms baselines on all precision metrics (except for large objects where it is comparable). Specifically, MOD -UV outperforms the nearest competitor (with ground-truth filtering )by 4.1 on Box AP and 3.4 on Mask AP, and is 4.5 higher on AP. Intriguingly, compared to a supervised Mask R-CNN trained on COCO [29], MOD -UV closes the gap in Mask AP from 4.3 to 0.8 points.
4.2.2 Generalization to Out-of-Domain Data.
We next take our detector trained on Waymo, and apply it out of the box on nuScenes, KITTI, and COCO.
Recall. As shown in Table 2 and 3, on nuScenes and KITTI, MOD -UV consistently outperforms prior arts across all AR metrics except for large objects, achieving a more than 1.5 improvement on AR0.5 over the nearest competitor on nuScenes. MOD -UV also shows large gains on small objects, improving AR by 2.4 on nuScenes and over 1.7 on KITTI.
Finally, we evaluate on COCO, which is in-domain for HASSOD and a big domain shift for MOD -UV that is trained on Waymo. Notably, MOD -UV maintains superiority on AR and AR at IoU, while being comparable to HASSOD on AR and AR.
Precision. This improvement is also seen in AP. On both nuScenes and KITTI, MOD -UV consistently outperforms baseline on all AP metrics except being comparable for APwith prior arts with ground-truth filtering. Specially, MOD -UV improves upon HASSOD∗(with ground truth filtering) on Box AP by 1.2 on nuScenes and 2.6 on KITTI, with notable improvements on AP by over 1.8 on nuScenes and 2.6 on KITTI. Even when compared on COCO which is the training set for HASSOD, MOD -UV outperforms prior arts on all metrics without ground-truth filtering except AP.
4.3 Qualitative Results
In Figure 3, we show qualitative examples of MOD -UV against CutLER and HASSOD after ground truth filtering, which we denote by CutLER∗ and HASSOD∗, respectively. In addition, we highlight the regions containing small objects with an additional zoom-in.
Without using any annotations, MOD -UV detects mobile objects accurately, especially recovering many more small and faraway objects compared to prior arts. In contrast, due to the reliance on image features from static images, both CutLER and HASSOD tend to group multiple objects into a single proposal (seen in the second row in Waymo).
Notably, MOD -UV reliably detects static and small mobile objects in the scene without excess amount of false positives. This improvement mostly originates from the proposed Moving2Mobile and Large2Small (see ablation below).
Beyond accurate detection and segmentation on Waymo, MOD -UV demonstrates impressive generalization when applied on nuScenes, KITTI, and COCO.
4.4 Ablation Study
We conduct ablation studies with MOD -UV‡ trained solely from Waymo [44] to understand the effects of each proposed component, including pseudo-label generation, static object discovery, small object discovery, and final round training.
4.4.1 Pseudo-Label Generation.
In Table 4, we measure the quality of the pseudo-labels in terms of AR for All, the Static, and the Moving instances.
In the top of Table 4, compared to using 2D contours for generate pseudo-labels, clustering pseudo 3D points from monocular depth estimations improves Moving AR by 9.4 and Moving AR by 2.5. This underlines the benefit in leveraging 3D information for partitioning moving instances. Nevertheless, the Static AR is notably smaller, with Static AR only up to 10% of Moving AR. Also, small and medium objects are almost entirely missed, with All AR at 0.0 and All AR at 0.5, compared to All AR at 10.0. These observations further verify the bias pointed out in Section 3.1, where static and small objects are mostly absent in the initial pseudo-labels generated from motion segmentation.
Furthermore, in the bottom half of Table 4, we demonstrate the effectiveness of self-training in the Moving2Mobile stage in recovering static objects from the initial pseudo-labels with varying number of training epochs. It is worth noting that regardless of convergence, self-training successfully improves Static AR, with improvements as high as 23.4 on Static AR. Interestingly, as the initial pseudo-labels contain mostly moving objects, additional training beyond 3 epochs shows clear degradation in performance (reducing Static AR by 5.0 while retaining Moving AR), as the trained detector overfits to moving instances by exploiting contextual priors. In addition to improvements on All AR, the Moving2Mobile stage is also able to slightly lift up All AR by the highest at 4.1. Nevertheless, with no improvements on AR, it is clear that the Moving2Mobile stage alone cannot alleviate the negative bias from motion segmentation.
4.4.2 Self-Training Pipeline.
In Table 5, we evaluate every combination of the 3-stage self-training pipeline with MOD -UV‡ to evaluate the effectiveness of each. Here, the Moving2Mobile stage is again shown to be essential for recovering static objects from the initial pseudo-labels. When Moving2Mobile is ablated, performance decrease across all combinations, with notable reductions in AR by 4.9 and AP by 3.9 when solely ablated from MOD -UV‡.
Additionally, when Large2Small is ablated, AR reduces by nearly 4 and AP by 3, underlining its importance for small object detection. Lastly, the final self-training round effectively learns the aggregated proposals from Large2Small, leading to an improved Box AR by 8.1 from arge-object detector trained at original scale and by 4.1 from mall-object detector trained at reduced scale.
5 Conclusion
We argue that motion is an important cue for unsupervised object detection, and propose the task of unsupervised mobile object detection. We propose a new training pipeline, MOD -UV, that bootstraps from motion segmentation but removes its bias by discovering static and small objects. MOD -UV achieves significant improvement over prior self-supervised detectors on multiple datasets.
Limitations and Societal Impact. Our work makes an assumption that all mobile objects would often move in the observed videos. Although MOD -UV can ideally learn and detect all mobile objects, but in practice, the learning-based framework can only learn and detect things that move in the videos. Being an unsupervised detection framework, our work does not include any negative social impacts beyond object detection itself.
Appendix 0.A Implementation Details
0.A.1 Mask Aggregation
When aggregating the predictions of “large object” detector, , and “small object” detector, , we found that contains mostly parts of large objects and small objects, while contains mostly large objects and groups of small objects. Intuitively, NMS is less-suited for this aggregation task due to the presence of object parts and groups.
Thus, we implement our aggregation as shown in Algo. 1. Here, we first filter out smaller overlapping masks in (e.g. ones that are covered by larger masks by more than filtFrac of 0.75) as the smaller objects should be found via instead. We also filter out larger overlapping masks in with the same filtFrac as the larger objects should be found via instead.
After directly matching the masks in and with a matchThrd of 0.5, if a subset of proposals in sufficiently covers a mask in (over a coverFrac of 0.5), then we consider the large proposal to likely be a group of instances and only keep the subset. Conversely, we consider the subset to likely be object parts and keep the large proposal.
As shown in Table 8, our aggregation approach improves upon NMS with a matchThrd of 0.5 by 1.8 in AR0.5 and 1.5 in AP.
Appendix 0.B Evaluation Datasets
0.B.1 Waymo Open Dataset
We evaluate performance on Waymo via all images in the val set of Waymo Open Dataset [44]. We obtain the instance-level object masks via the panoptic annotations [32], and treat the following object categories to be mobile: car, truck, bus, other_vehicle, bicycle, motorcycle, trailer, pedestrian, bicyclist, motorcyclist, bird, and ground_animal.
In total, there are 1,881 test images with 53,387 mobile instances labeled.
0.B.2 nuScenes Dataset
We evaluate performance on nuScenes via all FRONT camera images in the val set of nuImage [4]. We consider all object categories under the super-categories (animal,human, and vehicle) to be mobile. In total, there are 3,249 test images with 26,618 mobile instances labeled.
0.B.3 KITTI Dataset
We evaluate performance on KITTI via all images in the train set of KITTI 2D Detection Dataset [17]. We consider all object categories labeled in the dataset to be mobile. In total, there are 7,481 test images with 51,865 mobile instances labeled.
0.B.4 COCO Dataset
We evaluate performance on COCO via all images that contain street vehicles in the val2017 set [29]. We consider the following object categories to be mobile: car, truck, bus, bicycle, motorcycle.
In total, there are 870 test images with 3,319 mobile instances labeled. For reference, the original val2017 set contains 5,000 images with 36,781 labeled instances (regardless of mobility).
Method | Train | Init. | Box | Mask | ||||||||||||||
IoU=.50 | IoU=.50:.05:.95 | IoU=.50 | IoU=.50:.05:.95 | |||||||||||||||
AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | |||
Sup. Mask R-CNN [19] | 54.3 | 33.2 | 84.3 | 95.1 | 31.9 | 13.4 | 53.0 | 79.8 | 48.9 | 25.1 | 80.8 | 96.9 | 27.5 | 9.2 | 46.3 | 78.9 | ||
CutLER[52] | IN | IN | 20.9 | 3.1 | 33.1 | 87.5 | 11.7 | 1.3 | 17.3 | 54.1 | 20.0 | 2.5 | 30.8 | 86.8 | 10.7 | 0.9 | 14.9 | 52.7 |
HASSOD[6] | COCO | IN | 21.9 | 4.3 | 33.6 | 88.2 | 12.7 | 1.8 | 17.6 | 59.5 | 20.7 | 3.6 | 30.8 | 87.3 | 11.0 | 1.2 | 14.2 | 55.6 |
HASSOD†[6] | W | IN | 15.3 | 1.1 | 22.9 | 73.2 | 8.3 | 0.5 | 11.0 | 43.7 | 14.6 | 0.8 | 20.2 | 74.6 | 7.2 | 0.2 | 7.9 | 42.6 |
Ours | W | W | 40.0 | 23.0 | 56.1 | 92.5 | 17.5 | 8.4 | 25.7 | 46.4 | 35.4 | 19.0 | 49.2 | 89.0 | 14.6 | 6.8 | 21.2 | 40.4 |
Ours | W | IN | 39.9 | 21.3 | 59.9 | 92.2 | 19.3 | 8.6 | 28.5 | 54.1 | 35.8 | 17.6 | 53.2 | 90.0 | 16.4 | 7.2 | 23.8 | 47.3 |
AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | |||
Sup. Mask R-CNN [19] | 46.1 | 22.6 | 74.6 | 91.5 | 25.7 | 8.0 | 42.2 | 72.2 | 41.9 | 15.9 | 71.1 | 94.7 | 22.4 | 5.3 | 36.2 | 73.0 | ||
CutLER[52] | IN | IN | 8.8 | 1.1 | 7.5 | 55.7 | 5.0 | 0.5 | 3.9 | 32.0 | 9.1 | 0.1 | 7.1 | 58.5 | 5.2 | 0.0 | 3.4 | 34.6 |
HASSOD[6] | COCO | IN | 5.0 | 0.2 | 4.7 | 44.3 | 3.1 | 0.1 | 2.5 | 28.4 | 5.0 | 0.1 | 4.3 | 45.9 | 2.8 | 0.0 | 2.0 | 27.7 |
HASSOD†[6] | W | IN | 3.7 | 0.0 | 2.3 | 29.6 | 2.2 | 0.0 | 1.1 | 17.2 | 3.9 | 0.0 | 2.0 | 33.8 | 2.0 | 0.0 | 0.9 | 18.3 |
Ours | W | W | 26.1 | 12.3 | 36.5 | 75.0 | 10.9 | 4.2 | 16.2 | 32.0 | 25.0 | 11.0 | 34.8 | 75.2 | 9.5 | 3.7 | 14.1 | 28.7 |
Ours | W | IN | 26.3 | 12.3 | 37.9 | 76.5 | 12.6 | 4.9 | 17.9 | 41.0 | 25.1 | 10.5 | 35.9 | 77.5 | 11.1 | 4.5 | 15.6 | 36.3 |
CutLER∗[52] | IN | IN | 14.3 | 2.8 | 23.0 | 69.5 | 7.4 | 1.3 | 10.7 | 38.4 | 14.3 | 2.1 | 22.0 | 72.2 | 7.4 | 0.9 | 9.6 | 41.0 |
HASSOD∗[6] | COCO | IN | 15.3 | 3.0 | 23.3 | 68.6 | 8.5 | 1.4 | 10.9 | 42.6 | 15.0 | 2.3 | 21.3 | 70.6 | 7.7 | 1.0 | 8.9 | 41.3 |
HASSOD†∗[6] | W | IN | 9.3 | 1.3 | 15.2 | 46.4 | 4.7 | 0.6 | 6.5 | 25.2 | 9.6 | 0.3 | 13.4 | 52.7 | 4.5 | 0.1 | 4.9 | 27.0 |
Method | Train | Init. | Box | Mask | ||||||||||||||
IoU=.50 | IoU=.50:.05:.95 | IoU=.50 | IoU=.50:.05:.95 | |||||||||||||||
AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | |||
CutLER[52] | IN | IN | 28.9 | 8.6 | 43.9 | 88.2 | 16.0 | 3.8 | 23.4 | 56.1 | 27.8 | 7.5 | 41.7 | 87.6 | 14.3 | 3.0 | 20.4 | 53.0 |
HASSOD[6] | COCO | IN | 30.9 | 11.0 | 45.3 | 90.2 | 17.0 | 5.0 | 24.0 | 56.8 | 29.7 | 10.2 | 42.9 | 88.2 | 14.7 | 3.9 | 20.1 | 53.6 |
HASSOD†[6] | W | IN | 24.3 | 6.2 | 36.3 | 81.6 | 12.7 | 2.7 | 18.0 | 48.2 | 23.0 | 5.2 | 33.8 | 80.5 | 10.7 | 1.9 | 14.0 | 45.7 |
Ours | W | W | 42.1 | 24.3 | 57.8 | 85.5 | 17.3 | 8.7 | 24.2 | 39.8 | 36.4 | 23.5 | 46.4 | 70.7 | 13.9 | 8.2 | 18.1 | 29.7 |
Ours | W | IN | 48.9 | 31.8 | 65.0 | 87.8 | 21.9 | 12.0 | 29.8 | 48.2 | 42.3 | 27.4 | 54.8 | 78.8 | 18.3 | 10.7 | 24.0 | 39.1 |
AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | |||
CutLER[52] | IN | IN | 6.0 | 0.6 | 6.6 | 37.4 | 3.7 | 0.3 | 3.1 | 23.7 | 5.9 | 0.2 | 6.1 | 37.3 | 3.5 | 0.1 | 2.8 | 22.7 |
HASSOD[6] | COCO | IN | 3.9 | 0.3 | 4.4 | 33.6 | 2.2 | 0.1 | 2.1 | 20.5 | 3.8 | 0.2 | 4.1 | 33.1 | 2.0 | 0.1 | 1.8 | 19.5 |
HASSOD†[6] | W | IN | 3.6 | 0.1 | 3.5 | 26.2 | 2.2 | 0.0 | 1.7 | 15.2 | 3.6 | 0.1 | 2.5 | 27.0 | 1.8 | 0.0 | 0.9 | 14.6 |
Ours | W | W | 18.8 | 7.4 | 27.3 | 53.6 | 7.3 | 2.6 | 10.4 | 22.3 | 17.1 | 7.3 | 23.7 | 45.4 | 6.0 | 2.4 | 8.0 | 17.2 |
Ours | W | IN | 23.6 | 11.5 | 32.6 | 59.3 | 10.7 | 4.3 | 14.9 | 31.5 | 21.8 | 10.1 | 29.8 | 55.2 | 9.0 | 3.8 | 12.2 | 25.6 |
CutLER∗[52] | IN | IN | 15.6 | 5.6 | 26.8 | 65.7 | 8.2 | 2.3 | 12.6 | 38.4 | 15.2 | 4.7 | 25.2 | 65.4 | 7.7 | 1.7 | 11.4 | 37.0 |
HASSOD∗[6] | COCO | IN | 18.6 | 6.0 | 28.8 | 66.5 | 9.5 | 2.3 | 13.3 | 38.2 | 17.9 | 5.5 | 26.8 | 64.8 | 8.6 | 1.8 | 11.5 | 36.2 |
HASSOD†∗[6] | W | IN | 13.0 | 3.3 | 21.3 | 52.7 | 6.3 | 1.2 | 9.1 | 27.9 | 12.8 | 3.0 | 19.8 | 54.2 | 5.6 | 1.1 | 7.5 | 27.2 |
Mask Agg | Box AR | Box AP | ||||||||
Strategies | AR0.5 | AR | AR | AR | AR | AP50 | AP | AP | AP | AP |
NMS | 38.2 | 17.0 | 7.7 | 24.6 | 48.2 | 24.6 | 10.5 | 4.2 | 14.7 | 35.2 |
0.5/0.5 | 39.0 | 17.2 | 7.8 | 25.9 | 45.7 | 24.3 | 10.0 | 4.2 | 16.2 | 28.9 |
Ours | 40.0 | 17.5 | 8.4 | 25.7 | 46.4 | 26.1 | 10.9 | 4.2 | 16.2 | 32.0 |
Method | Train | Init. | KITTI [17] | COCO [29] | ||||||||||||||
IoU=.50 | IoU=.50:.05:.95 | IoU=.50 | IoU=.50:.05:.95 | |||||||||||||||
AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | AR | |||
CutLER[52] | IN | IN | 50.9 | 25.0 | 50.4 | 83.3 | 24.4 | 11.2 | 23.0 | 42.9 | 49.0 | 19.3 | 64.9 | 94.6 | 27.9 | 9.8 | 33.9 | 61.6 |
HASSOD[6] | COCO | IN | 52.4 | 29.6 | 50.5 | 83.9 | 26.4 | 13.4 | 23.9 | 47.1 | 51.2 | 24.1 | 66.6 | 91.7 | 27.9 | 12.5 | 36.2 | 51.7 |
HASSOD†[6] | W | IN | 49.9 | 33.2 | 48.5 | 73.3 | 23.7 | 15.0 | 22.1 | 37.4 | 48.8 | 24.5 | 59.4 | 90.5 | 26.1 | 12.5 | 31.2 | 50.7 |
Ours | W | W | 66.1 | 49.5 | 67.6 | 83.4 | 29.3 | 21.2 | 29.0 | 39.6 | 55.9 | 48.7 | 61.3 | 64.8 | 23.3 | 20.1 | 26.9 | 25.5 |
Ours | W | IN | 68.1 | 54.1 | 68.3 | 84.7 | 32.6 | 23.7 | 32.0 | 44.3 | 58.3 | 49.6 | 62.9 | 72.1 | 26.6 | 22.3 | 29.8 | 32.0 |
AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | AP | |||
CutLER[52] | IN | IN | 18.6 | 0.9 | 14.5 | 49.7 | 8.6 | 0.4 | 5.6 | 24.0 | 9.8 | 0.8 | 5.5 | 37.2 | 5.6 | 0.4 | 2.4 | 21.6 |
HASSOD[6] | COCO | IN | 14.3 | 0.7 | 11.4 | 55.7 | 7.2 | 0.3 | 5.1 | 29.5 | 3.1 | 0.3 | 4.0 | 16.8 | 1.4 | 0.1 | 1.9 | 7.6 |
HASSOD†[6] | W | IN | 16.7 | 1.4 | 16.2 | 42.5 | 7.5 | 0.6 | 6.8 | 19.3 | 4.7 | 0.6 | 5.4 | 21.0 | 2.7 | 0.3 | 3.1 | 10.0 |
Ours | W | W | 38.5 | 24.1 | 38.9 | 60.0 | 15.8 | 9.8 | 15.3 | 26.2 | 14.1 | 12.0 | 18.4 | 16.3 | 5.8 | 5.1 | 7.6 | 5.9 |
Ours | W | IN | 38.9 | 27.0 | 39.2 | 55.7 | 18.0 | 11.7 | 18.0 | 28.4 | 14.2 | 11.7 | 19.1 | 16.2 | 6.6 | 5.6 | 9.0 | 6.6 |
CutLER∗[52] | IN | IN | 28.4 | 9.2 | 26.1 | 61.7 | 12.3 | 3.5 | 9.9 | 28.8 | 24.3 | 9.9 | 27.6 | 70.3 | 12.8 | 4.3 | 12.1 | 40.1 |
HASSOD∗[6] | COCO | IN | 33.4 | 12.1 | 27.9 | 67.7 | 15.4 | 4.4 | 11.5 | 34.9 | 21.3 | 12.3 | 28.6 | 40.6 | 10.1 | 5.4 | 13.8 | 18.9 |
HASSOD†∗[6] | W | IN | 30.8 | 15.3 | 28.7 | 53.1 | 12.7 | 5.7 | 11.1 | 23.3 | 22.8 | 13.5 | 27.2 | 46.3 | 10.6 | 5.8 | 12.7 | 21.3 |
Appendix 0.C Additional Quantitative Results
0.C.1 Complete Metrics
0.C.2 Reproducibility
To ensure reproducibility, we repeat the experiment for MOD -UV‡ 3 times with randomly generated seeds and obtained a 95% confidence interval of 16.9 1.3 and 10.3 1.5 for Box AR and Box AP, respectively on Waymo Open. Please refer to Table 10 for standard deviations for each metric.
0.C.3 Mask Aggregation
In Table 8, we evaluate the effects of different mask aggregation techniques on the performance of the final detector. Notably, different strategies have minimal effects on Box AR and Box AP, which suggests the robustness of the mask aggregation step. Additionally, we do observe small performance gain for medium objects against the Non-Maximum Suppression (NMS) technique and small performance gain for large objects against the use of lowered confidence thresholds, “0.5/0.5” instead of “0.9/0.8”.
0.C.4 Backbone Pre-training.
In Table 11, we evaluate the effects of different backbone initializations on our proposed learning scheme. Since MOD -UV depends on multiple rounds of self-training, training detectors with a backbone initialized from scratch () reduces performance significantly. Interestingly, the use of any particular pre-training technique is less influential, as MOD -UV demonstrate similar performances.
Run | Box AR | Box AP | ||||||||
AR0.5 | AR | AR | AR | AR | AP50 | AP | AP | AP | AP | |
Standard Deviation | 1.3 | 0.8 | 0.8 | 1.1 | 3.5 | 0.8 | 0.9 | 0.3 | 1.3 | 4 |
Training | Init. Enc | Box AR | Box AP | ||||||||
Data | Data | AR0.5 | AR | AR | AR | AR | AP50 | AP | AP | AP | AP |
W | 29.9 | 13.4 | 5.9 | 18.5 | 40.7 | 16.0 | 7.1 | 2.6 | 10.5 | 25.9 | |
W | W | 40.0 | 17.5 | 8.4 | 25.7 | 46.4 | 26.1 | 10.9 | 4.2 | 16.2 | 32.0 |
W | IN | 39.9 | 19.3 | 8.6 | 28.5 | 54.1 | 26.3 | 12.6 | 4.9 | 17.9 | 41.0 |
W | IN(L) | 39.9 | 18.7 | 9.1 | 27.5 | 48.6 | 26.3 | 11.8 | 5.3 | 17.5 | 33.6 |

### v3

@@
-Learning to Detect and Segment Mobile Objects from Unlabeled Videos
+MOD-UV: Learning Mobile Object Detectors from Unlabeled Videos
 Abstract
@@
 Supervision in the form of bounding boxes for this task is extremely expensive.
-As such, prior work has looked at unsupervised object segmentation, but in the absence of annotated boxes, it is unclear how pixels must be grouped into objects and which objects are of interest.
+As such, prior work has looked at unsupervised instance detection and segmentation, but in the absence of annotated boxes, it is unclear how pixels must be grouped into objects and which objects are of interest.
 This results in over-/under-segmentation and irrelevant objects.
-Inspired both by the human visual system and by practical applications, we posit that the key missing cue is motion: objects of interest are typically mobile objects.
-We propose MOD -UV, a Mobile Object Detector learned from Unlabeled Videos only.
-We begin with pseudo-labels derived from motion segmentation, but introduce a novel training paradigm to progressively discover small objects and static-but-mobile objects that are missed by motion segmentation.
+Inspired by human visual system and practical applications, we posit that the key missing cue for unsupervised detection is motion: objects of interest are typically mobile objects that frequently move and their motions can specify separate instances.
+In this paper, we propose MOD-UV, a Mobile Object Detector learned from Unlabeled Videos only.
+We begin with instance pseudo-labels derived from motion segmentation, but introduce a novel training paradigm to progressively discover small objects and static-but-mobile objects that are missed by motion segmentation.
 As a result, though only learned from unlabeled videos, MOD-UV can detect and segment mobile objects from a single static image.
-Empirically, we achieve state-of-the-art performance in unsupervised mobile object detection on Waymo Open [44], nuScenes [4], and KITTI [17] Dataset without using any external data or supervised models.
-Code is publicly available at %definitionshttps://github.com/YihongSun/MOD-UV.
+Empirically, we achieve state-of-the-art performance in unsupervised mobile object detection on Waymo Open, nuScenes, and KITTI Datasets without using any external data or supervised models.
+Code is available at github.com/YihongSun/MOD-UV.
 Keywords:
@@
 The limited set of classes further runs the risk of missing important object categories (e.g.
-snowplows for self-driving cars).
+snowplows for self-driving applications).
 These concerns have motivated research into unsupervised object detection techniques that automatically discover objects from unlabeled data [54, 55, 6].
@@
 However, while promising, these approaches often produce many uninteresting and irrelevant “objects” in cluttered scenes (e.g.
-roads) and over- or under-segment objects of interest (e.g.
-multiple detections partitioning a large bus and a single detection grouping a row of parked cars).
-These failures shouldn’t be surprising: after all, how can a completely unsupervised feature representation encode that one assortment of windows, doors and wheels belongs together as an object of interest, but is a separate object from other nearby windows, doors and wheels?
-In this paper, we argue that a key missing cue is motion.
-As humans, we know to group the windows, doors and wheels together because we have seen this assortment move together, also known as Gestalt principle of common fate [34].
-Having seen this assortment move together multiple times in the past, we can now group these correctly into an object even when the object (a car) is static.
-Indeed, motion-based grouping is one of the first forms of grouping to appear developmentally in human infants [43] and can potentially bootstrap other grouping cues [33].
-Motion is important not just perceptually, but also practically: Robots performing navigation tasks must plan their trajectories carefully around objects that can move.
-Identifying mobile objects is also important in manipulation tasks because it provides a view of the available affordances.
-With these motivations in mind, we propose an unsupervised approach to build a detector for mobile objects as shown in Figure 1.
-Our approach trains on unlabeled videos alone and produces a mobile object detector that can run on static frames.
-We build upon recent advances in unsupervised learning of depth and motion segmentation from unlabeled video to automatically produce pseudo training labels without supervision.
-However, such unsupervised motion segmentation poses two key challenges.
-One, it only identifies moving objects and does not include objects (e.g.
+buildings and roads) and over- or under-segment objects of interest (e.g.
+multiple detections partitioning a large bus or a single detection grouping a row of parked cars).
+These failures shouldn’t be surprising: after all, how can a completely unsupervised feature representation encode which assortment of windows, doors and wheels belongs together as an object, and which objects are of interest?
+In this paper, we argue that a key missing cue for addressing the aforementioned issues in unsupervised instance detection is motion.
+In a practical sense, objects of interest are commonly mobile objects that frequently move.
+For example, robots performing navigation tasks must plan their trajectories carefully around objects that can move of their own volition.
+Thus, we argue that if we see similar groups of pixels frequently move of their own volition in unlabeled videos (e.g.
+vehicles and pedestrians in driving videos), this is sufficient information for building a mobile object detector that can detect such instances in static frames.
+The importance of motion as a perceptual cue (the Gestalt principle of common fate) is well known [35].
+Indeed, motion-based grouping is one of the first forms of grouping to appear developmentally in human infants [45] and can bootstrap other grouping cues [34].
+There is also some prior work on using motion-based grouping in computer vision to produce (pseudo) ground-truth for feature learning [36] and discovering isolated, salient objects [11, 10].
+However, there is a big-gap between motion segmentation and the kind of ground-truth we need for training a full-fledged instance-level object detector.
+First, motion segmentation produces a binary segmentation; this must be resolved into individual instances.
+Second, it only identifies moving objects and does not include objects (e.g.
 parked cars) that are static but mobile.
-Two, it only identifies nearby objects, since the pixel motion of faraway objects is too subtle to discern.
-We propose a new training scheme to address these challenges, resulting in a final mobile object detector that detects more mobile objects than the initial motion segmentation.
+Finally, motion segmentation only identifies nearby objects, since the pixel motion of faraway objects is too subtle to discern.
+Thus motion segmentation alone will still under-segment and miss many mobile objects: a problem for building object detectors.
+Here we propose a new training scheme to address these challenges.
+Our approach (MOD-UV, a Mobile Object Detector learned from Unlabeled Videos only; Figure 1) trains on unlabeled videos alone and produces a mobile object detector that can run on static frames.
+We first generate pseudo training labels from motion segmentation estimated by our prior unsupervised framework Dynamo-Depth [47].
+We then propose a new training scheme to address the challenges above, resulting in a final mobile object detector that detects more mobile objects than the initial motion segmentation.
 We test MOD-UV on self-driving scenes but evaluate on a variety of datasets.
@@
 1.
-We argue that motion is an important cue for learning what makes an object.
+We argue that motion as a cue is sufficient for unsupervised training of instance-level object detectors.
 2.
-We propose a new training scheme that trains on unlabeled video to produce a mobile object detector that can run on static images.
+We propose a new training scheme that trains on unlabeled videos to produce a mobile object detector that can run on static images.
 3.
@@
 2.0.1 Unsupervised Object Detection/Discovery from Images.
-Learning to identify and localize objects from unlabeled images is a challenging task, as object information must be obtained without any explicit human annotations.
-A long line of work seeks to discover prominent objects in large image collections.
-Early work used probabilistic Hough transforms to match proposals across images[9].
-This approach was improved upon by framing the problem as an optimization problem over a graph where nodes are object proposals and edges are visual similarities[48] and scaled up to larger collections [49, 50].
-However, this approach is fundamentally limited by the quality of the object proposals.
+Learning to identify and localize objects from unlabeled images is a challenging task, since object information must be obtained without any explicit human annotations.
+A long line of work seeks to discover prominent objects in large image collections [9, 51, 52, 53].
+However, these approaches are fundamentally limited by the quality of the object proposals.
 More recently, Locatello et al.
-[30] and DINOSAUR [36] consider object discovery as object-centric learning [18] and leverage Slot Attention to decompose a complex scene into independent objects.
-However the reconstruction objective is difficult to scale and is likely to discover irrelevant objects as well.
+[30] and DINOSAUR [38] consider object discovery as object-centric learning [18] and decompose a complex scene into independent objects.
+Nevertheless, the reconstruction objective is difficult to scale and likely to discover irrelevant patches as well.
 More recent work has relied on the fact that bottom-up segmentation algorithms when applied to self-supervised pretrained representations yield good object proposals.
-LOST [39], FOUND [40] and FreeSOLO [51] generate pseudo mask labels from DINO [7] features to train their downstream detectors.
-Similarly, MaskDistill [47] distills a single class-agnostic initial mask from the affinity graph produced by DINO [7].
-TokenCut [53] and CutLER [52] use Normalized Cuts [37] instead to generate pseudo-labels.
-HASSOD [6] leverages hierarchical adaptive clustering to generate pseudo-labels, which improves the detection of small objects and object parts.
+Specifically, pseudo mask labels can be generated from DINO [7] features to train downstream object detectors [41, 42, 54].
+MaskDistill [50] extends upon this by distilling from affinity graph produced by DINO [7] features, while TokenCut [56] and CutLER [55] use Normalized Cuts [39].
+HASSOD [6] leverages hierarchical adaptive clustering, which improves the detection of small objects and object parts.
 This line of work now produces detectors that can run on static images, similar to our work.
 However, the detected objects can often be irrelevant (e.g.
-buildings and road) or over- or under-segment objects of interest.
+buildings and road) or over-/under-segment objects of interest.
 In contrast, our proposed MOD-UV discovers and detects a more meaningful and practical set of mobile objects instead, and can learn from their apparent motion in unlabeled videos only without relying on any additional datasets.
 2.0.2 Unsupervised Object Detection/Discovery from 3D.
-In addition to unlabeled images, 3D information can also be leveraged for discovering objects.
+In addition to unlabeled images, 3D information is also useful for discovering objects.
 Herbst et al.
-[21] reconstructs the same scene from multiple traversals via RGB-D inputs and identify discrepancies as discovered mobile objects.
-MODEST [57] extends upon this and builds an unsupervised 3D detector from traversals of LiDAR scans.
+[21] and MODEST [62] discover non-persistent objects via multiple traversals with 3D sensors.
 Garcia et al.
-[16] discovers salient objects by late-fusing color and depth segmentation from RGB-D inputs.
-Tian et al.
-[46] generates candidate segments via 3D point cloud segmentation from LiDAR recordings.
+[16] discovers salient objects by late-fusing color and depth segmentation from RGB-D inputs, while Tian et al.
+[49] generates candidate segments from LiDAR 3D point clouds.
 In comparison, MOD-UV does not require any additional sensors or modalities beyond unlabeled videos, which allows our method to work in more general settings.
-2.0.3 Unsupervised Video Segmentation.
-Another class of related work segments out the moving objects from a video without annotations.
-Here the input is a video and the output is a binary segmentation of the moving objects.
-Kwak et al.
-[25], Xie et al.
-[54], Lamdouar et al.
-[26], Yang et al.
-[56], Zhang et al.
-[58], Zhou et al.
-[59], and FODVid [41] perform this task by leveraging optical flow information from an input video.
-Additionally, Lian et al.
-[28] improves upon cases of articulated/deformable objects and shadow/reflections by relaxing the common fate assumption.
-Another line of work uses a reconstruction objective to identify the moving object.
+2.0.3 Unsupervised Object Detection/Discovery from Videos.
+Inspired by Gestalt principle of common fate [35], another class of related work discovers objects via their apparent motions observed in videos [59, 61, 31].
+By leveraging optical flow information from an input video, a binary segmentation of the moving objects can be extracted [36, 25, 57, 26, 60, 63, 64, 43].
+Lian et al.
+[28] proposes further improvements for cases of articulated/deformable objects and shadow/reflections by relaxing the common fate assumption.
+Another line of work uses a reconstruction objective to identify the moving object [48, 1].
 Du et al.
-[13] models explicit physical dynamics and reconstructs 3D object geometry by exploiting motion cues in videos.
+[13] models explicit object geometry and physical dynamics by exploiting motion cues.
 Bao et al.
-[2] improves training for object-centric representation via an additional motion segmentation regularization, while SAVi++ [14] incorporates depth values from LiDAR sensor when training an object-centric video model.
-Aydemir et al.
-[1] extends DINOSAUR [36] to video learning while proposing a masking strategy for tokens in the feature space for efficiency and regularization.
+[2] improves training for object-centric representation via an additional motion segmentation regularization, while SAVi++ [14] incorporates LiDAR data when training an object-centric video model.
 Unlike MOD-UV, these approaches do not build a static image detector.
 However, the output segmentation can be used as an initialization for our approach.
-Closer to our work, Croitoru et al.
+Closer to our work, Pathak et al.
+[36], Croitoru et al.
 [11] and Choudhury et al.
-[10] train a single-frame binary segmentation network on frames of a video as input and leverage object motion as supervision.
+[10] train a single-frame binary segmentation network on video frames as input and leverage object motion as supervision.
 Furthermore, LOCATE [44] applies graph-cut to obtain binary motion mask from DINO [7] and optical flow feature similarities, which in turn is treated as pseudo-labels for bootstrapped self-training of a downstream segmentation network.
 However, these techniques can only detect a single salient object per frame.
-In contrast, MOD -UV generalizes to multi-object detection during inference beyond single-object saliency detection.
+In contrast, MOD-UV generalizes to multi-object detection beyond single-object saliency detection.
 3 Method
-Problem setup: We assume an uncurated collection of unlabeled videos as input.
+3.0.1 Problem setup:
+We assume an uncurated collection of unlabeled videos as input.
 In particular, we assume that these videos are obtained by an embodied agent observing, and optionally acting in the world.
@@
 3.1 Initialization with Unsupervised Motion Segmentation
-A key insights in MOD -UV is that if an object can move, it is likely that it does move many times in the collected data.
+A key insight in MOD-UV is that if an object can move, it is likely that it does move many times in the collected data.
 Thus, we start by identifying moving objects in the videos; they can be initial seeds for learning about mobile objects.
@@
 3.2.1 Moving2Mobile: Learning to Detect Static Objects.
-As shown in the left of Figure 2, the initial pseudo-labels , while having high precision, fail to capture the large static objects, e.g.
+On the left of Figure 2, the initial pseudo-labels , while having high precision, fail to capture the large static objects, e.g.
 the black sedan in the bottom.
@@
 Upon convergence, we have a Large-object detector trained at original scale and another Small-object detector trained at a reduced scale.
-After aggregating their predictions and resolving conflicting proposals222Details on the mask aggregation are found in Appendix 0.A.1., we obtain the final pseudo-labels, .
+After aggregating their predictions and resolving conflicting proposals, we obtain the final pseudo-labels, .
 We note that separating out large and small object detectors in this way has been explored in supervised face detection [22].
@@
 3.3 Implementation details
-We follow the official code release by Dynamo-Depth[45] and train the system on Waymo Open [44] only to obtain the motion segmentation network and monocular depth estimator.
-In the initial pseudo-label generation stage, we binarize the estimated motion mask via a threshold of and cluster the pseudo 3D points via DBSCAN [15] using a 10-by-10 local pixel neighborhood connectivity, Eps at 0.03, and required number of samples for core point at 100.
+We follow the official code release by Dynamo-Depth[47] and train the system on Waymo Open [46].
+During initial pseudo-label generation, we binarize the estimated motion mask via a threshold of and cluster the pseudo 3D points via DBSCAN [15] using a 10-by-10 local pixel neighborhood connectivity.
 We adopt Mask R-CNN [19] with a ResNet-50[20] backbone as the detector architecture.
@@
 Since early-stopping is applied, we adopt a lower confidence threshold of to compute the next round pseudo-labels .
-During Large2Small, we train both the large and small detectors for 20 epochs, with fixed scale jittering at and , respectively.
+During Large2Small, we train both the large and small detectors for 20 epochs, with fixed scaling at and , respectively.
 As both are trained to convergence, we adopt a higher confidence threshold of and , respectively, to compute the next round pseudo-labels with aggregation.
@@
 4.1 Experimental Setup
-4.1.1 Datasets.
+Datasets.
 For evaluation, we focus our attention to self-driving datasets since uncurated, unlabeled video data is available and detecting mobile objects is of interest for autonomous vehicles.
@@
 We test on 3,249 front-camera images (average of 8.2 mobile instances per image) from the nuImage validation set for nuScene and 7,481 images (average of 6.9 mobile instances per image) in the 2D Detection training set for KITTI.
-For COCO, we evaluate on 870 images (average of 3.8 mobile instances per image) in COCO val 2017 that contain ground vehicles.333Additional dataset information is found in Appendix 0.B.
-4.1.2 Baselines.
+For COCO, we evaluate on 870 images (average of 3.8 mobile instances per image) in COCO val 2017 that contain ground vehicles.
+Baselines.
 For the task of unsupervised mobile object detection, there is no directly comparable baselines to the best of our knowledge.
@@
 We evaluate the official checkpoint.
+Furthermore, we consider an additional baseline where is directly predicted via CutLER, which effectively ablates the use of motion cues, as denoted by CutLERL2S.
 HASSOD [6] is a follow-up to CutLER.
@@
 As before, we use its official released checkpoint.
-Since both MOD -UV‡ and MOD -UV are trained on Waymo [44], we also consider a version of HASSOD solely trained on Waymo Open [44], which we denote as HASSOD†.
+Since MOD-UV‡ is trained on Waymo Open [46], we also consider a version of HASSOD solely trained on Waymo, which we denote as HASSOD†.
 CutLER[55] and HASSOD[6] are trained to detect all objects in an image regardless of their ability to move.
@@
 We therefore also consider an “oracle” version of these baselines where we additionally remove any CutLER and HASSOD predictions that overlap by less than 0.1 in IoU with the ground truth instances.
-These oracles, namely CutLER∗, HASSOD∗, and HASSOD†∗, are grayed out in the tables to indicate additional ground-truth-based filtering.
+These oracles, namely CutLER∗, CutLERL2S∗, HASSOD∗, and HASSOD†∗, are grayed out in the tables to indicate additional ground-truth-based filtering.
 We also consider a fully-supervised Mask R-CNN (Sup. Mask R-CNN) trained on COCO [29] for an oracle comparison, marked in gray.
-4.1.3 Metrics.
+Metrics.
 We evaluate both Average Recall (AR and AR) and Average Precision (AP and AP).
-Since the task is unsupervised and no semantic information is given during training, we follow prior arts [52, 6] and evaluate class-agnostic AR and AP by treating all predicted and ground truth instances as a single class of “foreground” or “mobile” objects.444Complete benchmarks, repeated training performance, and additional ablation studies are found in Appendix 0.C.
+Since the task is unsupervised and no semantic information is given during training, we follow prior arts [55, 6] and evaluate class-agnostic AR and AP by treating all predicted and ground truth instances as a single class of “foreground” or “mobile” objects.
 Method | Train | Init. | Box | Mask | ||||||||
@@
 CutLER[55] | IN | IN | 20.9 | 11.7 | 1.3 | 17.3 | 54.1 | 20.0 | 10.7 | 0.9 | 14.9 | 52.7 |
+CutLERL2S[55] | IN+W | IN | 29.3 | 15.0 | 4.6 | 24.3 | 48.2 | 28.1 | 14.4 | 4.2 | 22.9 | 47.8 |
 HASSOD[6] | COCO | IN | 21.9 | 12.7 | 1.8 | 17.6 | 59.5 | 20.7 | 11.0 | 1.2 | 14.2 | 55.6 |
@@
 CutLER[55] | IN | IN | 8.8 | 5.0 | 0.5 | 3.9 | 32.0 | 9.1 | 5.2 | 0.0 | 3.4 | 34.6 |
+CutLERL2S[55] | IN+W | IN | 9.6 | 4.3 | 0.9 | 9.6 | 16.9 | 9.6 | 4.4 | 0.8 | 9.8 | 17.7 |
 HASSOD[6] | COCO | IN | 5.0 | 3.1 | 0.1 | 2.5 | 28.4 | 5.0 | 2.8 | 0.0 | 2.0 | 27.7 |
@@
 CutLER∗[55] | IN | IN | 14.3 | 7.4 | 1.3 | 10.7 | 38.4 | 14.3 | 7.4 | 0.9 | 9.6 | 41.0 |
+CutLERL2S∗[55] | IN+W | IN | 23.2 | 10.1 | 3.3 | 16.4 | 32.6 | 22.8 | 10.2 | 3.1 | 16.3 | 34.1 |
 HASSOD∗[6] | COCO | IN | 15.3 | 8.5 | 1.4 | 10.9 | 42.6 | 15.0 | 7.7 | 1.0 | 8.9 | 41.3 |
@@
 This suggests that we detect significantly more objects than prior work, but their localization can be improved.
-Even so, on overall AR we still show a -point improvement.
-Additionally, although HASSOD† is trained in-domain on Waymo, it underperforms HASSOD on Waymo, which we suspect is due to the uncurated nature of self-driving scenes.
+Even so, we still show a -point improvement on overall AR.
+We also found HASSOD to underperform when trained solely on Waymo, which we suspect is due to the uncurated nature of self-driving scenes.
 Precision.
@@
 On Waymo, MOD-UV significantly outperforms prior arts across all precision metrics.
-Even with ground-truth filtering (in gray), MOD -UV still consistently outperforms baselines on all precision metrics (except for large objects where it is comparable).
+Even with ground-truth filtering (in gray), MOD-UV still consistently outperforms baselines on all precision metrics (except for larger objects where it is comparable).
 Specifically, MOD-UV outperforms the nearest competitor (with ground-truth filtering) by 4.1 on Box AP and 3.4 on Mask AP, and is 4.5 higher on AP.
@@
 Recall.
-As shown in Table 2 and 3, on nuScenes and KITTI, MOD -UV consistently outperforms prior arts across all AR metrics except for large objects, achieving a more than 1.5 improvement on AR0.5 over the nearest competitor on nuScenes.
+As shown in Table 2 and Table 3, on nuScenes and KITTI, MOD-UV consistently outperforms prior arts across all AR metrics except for large objects, achieving a more than 1.5 improvement on AR0.5 over the nearest competitor on nuScenes.
 MOD-UV also shows large gains on small objects, improving AR by 2.4 on nuScenes and over 1.7 on KITTI.
-Finally, we evaluate on COCO, which is in-domain for HASSOD and a big domain shift for MOD -UV that is trained on Waymo.
-Notably, MOD -UV maintains superiority on AR and AR at IoU, while being comparable to HASSOD on AR and AR.
+Finally, we evaluate on COCO, which is in-domain for HASSOD and a big domain shift for MOD-UV. Notably, MOD-UV maintains superiority on AR and AR at IoU, while being comparable to HASSOD on AR and AR.
 Precision.
@@
 Specially, MOD-UV improves upon HASSOD∗(with ground truth filtering) on Box AP by 1.2 on nuScenes and 2.6 on KITTI, with notable improvements on AP by over 1.8 on nuScenes and 2.6 on KITTI.
-Even when compared on COCO which is the training set for HASSOD, MOD -UV outperforms prior arts on all metrics without ground-truth filtering except AP.
+Even on COCO, MOD-UV outperforms prior arts on all metrics without ground-truth filtering except AP.
 4.3 Qualitative Results
@@
 We conduct ablation studies with MOD-UV‡ trained solely from Waymo [46] to understand the effects of each proposed component, including pseudo-label generation, static object discovery, small object discovery, and final round training.
+4.4.1 Motion Cues for Initial Pseudo-Labels.
+As shown in Table 1, there is a consistent AR improvement for small and medium objects in CutLERL2S over CutLER.
+This highlights the effectiveness of our Large2Small strategy in improving detector performance on small objects, which is a challenge for all unsupervised detectors/object discovery techniques because of the limited signal on small objects.
+Despite of this gain, MOD-UV still outperforms CutLERL2S because motion offers a stronger cue to separate small objects that appear close to each other in pixel space, as shown in Figure 3.
 4.4.2 Pseudo-Label Generation.
@@
 MOD-UV achieves significant improvement over prior self-supervised detectors on multiple datasets.
-Limitations and Societal Impact.
-Our work makes an assumption that all mobile objects would often move in the observed videos.
-Although MOD -UV can ideally learn and detect all mobile objects, but in practice, the learning-based framework can only learn and detect things that move in the videos.
-Being an unsupervised detection framework, our work does not include any negative social impacts beyond object detection itself.
+Limitations.
+Our work makes an assumption that all mobile objects would often move in the given unlabeled video dataset.
+Although MOD-UV can ideally learn and detect all mobile objects, in practice, the learning-based framework can only learn and detect things that frequently move in the videos.
+That said, for general applications where the autonomous agents can manipulate their surroundings, the agent can still learn to detect rarely moving objects by interacting with their environment, e.g.
+poking at static objects within reach.
+Societal Impact. Being an unsupervised detection framework, our work does not include any negative social impacts beyond object detection itself.
+Acknowledgement
+This research is based upon work supported in part by the National Science Foundation (IIS-2144117 and IIS-2107161). Yihong Sun is supported by an NSF graduate research fellowship.
 Appendix 0.A Implementation Details
@@
 CutLER[55] | IN | IN | 20.9 | 3.1 | 33.1 | 87.5 | 11.7 | 1.3 | 17.3 | 54.1 | 20.0 | 2.5 | 30.8 | 86.8 | 10.7 | 0.9 | 14.9 | 52.7 |
+CutLERL2S[55] | IN+W | IN | 29.3 | 10.7 | 47.9 | 85.1 | 15.0 | 4.6 | 24.3 | 48.2 | 28.1 | 9.6 | 45.3 | 84.7 | 14.4 | 4.2 | 22.9 | 47.8 |
 HASSOD[6] | COCO | IN | 21.9 | 4.3 | 33.6 | 88.2 | 12.7 | 1.8 | 17.6 | 59.5 | 20.7 | 3.6 | 30.8 | 87.3 | 11.0 | 1.2 | 14.2 | 55.6 |
@@
 CutLER[55] | IN | IN | 8.8 | 1.1 | 7.5 | 55.7 | 5.0 | 0.5 | 3.9 | 32.0 | 9.1 | 0.1 | 7.1 | 58.5 | 5.2 | 0.0 | 3.4 | 34.6 |
+CutLERL2S[55] | IN+W | IN | 9.6 | 2.0 | 20.9 | 35.6 | 4.3 | 0.9 | 9.6 | 16.9 | 9.6 | 1.8 | 20.8 | 36.3 | 4.4 | 0.8 | 9.8 | 17.7 |
 HASSOD[6] | COCO | IN | 5.0 | 0.2 | 4.7 | 44.3 | 3.1 | 0.1 | 2.5 | 28.4 | 5.0 | 0.1 | 4.3 | 45.9 | 2.8 | 0.0 | 2.0 | 27.7 |
@@
 CutLER∗[55] | IN | IN | 14.3 | 2.8 | 23.0 | 69.5 | 7.4 | 1.3 | 10.7 | 38.4 | 14.3 | 2.1 | 22.0 | 72.2 | 7.4 | 0.9 | 9.6 | 41.0 |
+CutLERL2S∗[55] | IN+W | IN | 23.2 | 8.5 | 37.9 | 68.4 | 10.1 | 3.3 | 16.4 | 32.6 | 22.8 | 7.7 | 37.0 | 69.7 | 10.2 | 3.1 | 16.3 | 34.1 |
 HASSOD∗[6] | COCO | IN | 15.3 | 3.0 | 23.3 | 68.6 | 8.5 | 1.4 | 10.9 | 42.6 | 15.0 | 2.3 | 21.3 | 70.6 | 7.7 | 1.0 | 8.9 | 41.3 |
@@
 0.C.1 Complete Metrics
-0.C.2 Reproducibility
-To ensure reproducibility, we repeat the experiment for MOD -UV‡ 3 times with randomly generated seeds and obtained a 95% confidence interval of 16.9 1.3 and 10.3 1.5 for Box AR and Box AP, respectively on Waymo Open.
-Please refer to Table 10 for standard deviations for each metric.
-0.C.3 Mask Aggregation
+For additional performance metrics on Waymo Open [46], please refer to Table 6 that corresponds to Table 1.
+Also, nuScenes [4] performances can be found in Table 7 (corresponding to Table 2).
+Finally, KITTI [17] and COCO [29] results are found in Table 9 (corresponding to Table 3).
+Interestingly, we found that motion segmentation trained from a reconstruction objective [47] contains specific biases, such as the inclusion of object shadows and missing object parts due to smooth object regions.
+This results in localization errors and reduced performances for large objects for IoU>0.75.
+As shown in the Table 6, MOD-UV outperforms all prior art even for large objects for IoU=0.5 on Waymo Open.
+0.C.2 Mask Aggregation
 In Table 8, we evaluate the effects of different mask aggregation techniques on the performance of the final detector.
@@
 Additionally, we do observe small performance gain for medium objects against the Non-Maximum Suppression (NMS) technique and small performance gain for large objects against the use of lowered confidence thresholds, “0.5/0.5” instead of “0.9/0.8”.
-0.C.4 Backbone Pre-training.
+0.C.3 Backbone Pre-training.
 In Table 10, we evaluate the effects of different backbone initializations on our proposed learning scheme.
@@
 Interestingly, the use of any particular pre-training technique is less influential, as MOD-UV demonstrate similar performances.
-Run | Box AR | Box AP | ||||||||
-AR0.5 | AR | AR | AR | AR | AP50 | AP | AP | AP | AP | |
-Standard Deviation | 1.3 | 0.8 | 0.8 | 1.1 | 3.5 | 0.8 | 0.9 | 0.3 | 1.3 | 4 |
 Training | Init. Enc | Box AR | Box AP | ||||||||
@@
 W | IN(L) | 39.9 | 18.7 | 9.1 | 27.5 | 48.6 | 26.3 | 11.8 | 5.3 | 17.5 | 33.6 |
+Run | Box AR | Box AP | ||||||||
+AR0.5 | AR | AR | AR | AR | AP50 | AP | AP | AP | AP | |
+Standard Deviation | 1.3 | 0.8 | 0.8 | 1.1 | 3.5 | 0.8 | 0.9 | 0.3 | 1.3 | 4 |
+0.C.4 Reproducibility and Hyperparameter Selection
+To ensure reproducibility, we repeat the experiment for MOD-UV‡ 3 times with randomly generated seeds and obtained a 95% confidence interval of 16.9 1.3 and 10.3 1.5 for Box AR and Box AP, respectively on Waymo Open.
+Please refer to Table 11 for standard deviations for each metric.
+The number of epochs for Moving2Mobile, scale jittering rates, and confidence thresholds for self-training were found on a small Waymo validation set during the development of our paper, while the test set is held-out entirely.
+The rest of the hyperparameters, including number of epochs, learning rates, and decay, were set arbitrarily and not tuned since training converged.